### v1

# Compositional Reasoning with Transformers, RNNs, and Chain of Thought

Abstract
We study and compare the expressive power of transformers, RNNs, and transformers with chain of thought tokens on a simple and natural class of problems we term Compositional Reasoning Questions (CRQ). This family captures problems like evaluating Boolean formulas and multi-step word problems. Assuming standard hardness assumptions from circuit complexity and communication complexity, we prove that none of these three architectures is capable of solving CRQs unless some hyperparameter (depth, embedding dimension, and number of chain of thought tokens, respectively) grows with the size of the input. We also provide a construction for each architecture that solves CRQs. For transformers, our construction uses depth that is logarithmic in the problem size. For RNNs, logarithmic embedding dimension is necessary and sufficient, so long as the inputs are provided in a certain order. (Otherwise, a linear dimension is necessary). For transformers with chain of thought, our construction uses $n$ CoT tokens. These results show that, while CRQs are inherently hard, there are several different ways for language models to overcome this hardness. Even for a single class of problems, each architecture has strengths and weaknesses, and none is strictly better than the others.
1 Introduction
Large language models (Touvron et al., 2023; Anil et al., 2023; Achiam et al., 2023) are increasingly used to perform logical reasoning and other problems that require algorithmic thinking. To understand the power and limitations of these models, it is essential to determine what kinds of computational problems they are capable of solving. To this end, a long line of theoretical work has studied the expressive power of various language modeling architectures using simple tasks like copying strings, recognizing formal languages, and determining if a graph is connected (Jelassi et al., 2024; Sanford et al., 2024b; Strobl et al., 2024). Such studies typically provide constructions or prove impossibility results for a particular architecture or paradigm—such as recurrent neural networks, transformers, and transformers with test-time scaling (“chain of thought”)—following the progress of the field.
In this paper, we take a broader view. We compare the abilities of several different language modeling architectures to solve a class of problems, which we call Compositional Reasoning Questions (CRQs). Our findings prove that there are unavoidable trade-offs between the architectures we consider. No architecture is capable of solving CRQs without some strong dependence on the size of the problem. However, each architecture allows us to limit the use of some computational resource at the expense of the others (see Table 1). Thus, the CRQ task provides a crisp and formal way to characterize the essential differences between these architectures.
Our Compositional Reasoning Questions are inspired by the complex, multi-step questions for which LLMs are often used in practice. For example, consider the question: Amongst the record title winners of each Grand Slam, who is the player with the fewest overall titles? Answering this question requires solving several sub-questions. While some of them can be solved in parallel (i.e., finding the record holders for each Grand Slam), the overall question cannot be answered without first gathering the answers to the sub-questions. It shares this structure with arithmetic and boolean expressions like $(6+2)\cdot(4-5)$ and is also similar to word problems and logic puzzles. Problems like these can be thought of as having an underlying tree structure. The leaves are the possible answers, each non-leaf node represents a sub-question, and the overall answer is the answer to the root node, see Figure 1 for an example. This type of compositional reasoning is a fundamental challenge for LLMs.
Table 1: Comparison between our methods for solving CRQs on $n$ nodes. Each architecture minimizes one kind of resource at the expense of the others. For ease of comparison, we assume the depth of the CRQ tree is $\log n$. By parallel runtime, we mean the runtime using an unlimited number of parallel processors.
Architecture Num. Parameters Runtime Parallel Runtime
Deep transformer (Section 4) $O(\log^{2}n)$ $O(n^{2}\,\log^{2}n)$ $O(\log n)$
RNN (Section 5) $O(\log n)$ $O(n\log n)$ $O(n)$
Chain of Thought (Section 6) $O(1)$ $O(n^{2})$ $O(n)$
We study the abilities of deep transformers, recurrent neural networks, and shallow transformers with chain of thought to solve these problems. Our main contributions are as follows:
  1. 1.
    In Section 3.1, we present Compositional Reasoning Questions, a simple, formal framework based on semantic similarity for studying LLM reasoning on arbitrary tree structures.
  2. 2.
    In Section 4, we prove that transformers with constant depth cannot solve arbitrary CRQs (Theorem 4.3), but transformers with depth $L$ can solve all CRQs of depth up to $L$ (Theorem 4.1).
  3. 3.
    In Section 5, we prove that RNNs with constant hidden dimension cannot solve arbitrary CRQs (Theorem 5.5), but RNNs with $O(\log n)$ hidden dimension and constant depth can solve all CRQs of size $n$ (Theorem 5.4). This ability depends on the inputs being arranged in a particular order (Algorithm 1); if they are ordered adversarially, RNNs require $O(n)$ hidden dimension (Theorem 5.2).
  4. 4.
    In Section 6, we prove that transformers augmented with $O(\log n)$ CoT tokens cannot solve CRQs of size $n$, but transformers augmented with $O(n)$ CoT tokens can (Theorem 6.1).
These results demonstrate the trade-offs between different models (Table 1). While deep transformers are highly parallelizable, they require $O(\log n)$ depth in the worst case, so the model size must depend (albeit mildly) on the problem size. Likewise, RNNs use little compute but must grow to handle larger problems. Chain of thought allows a single, constant-size model to handle any CRQ, but it runs slowly and is not parallelizable. Overall, our work reveals a nuanced complexity landscape for an important class of reasoning problems.
$\begin{bmatrix}1\\ 1\end{bmatrix}$$\begin{bmatrix}1\\ 0\end{bmatrix}$$\begin{bmatrix}7\\ 6\end{bmatrix}$$\begin{bmatrix}2\\ 3\end{bmatrix}$$\begin{bmatrix}4\\ 6\end{bmatrix}$$\begin{bmatrix}0\\ -1\end{bmatrix}$$\begin{bmatrix}2\\ 8\end{bmatrix}$$\begin{bmatrix}3\\ 4\end{bmatrix}$$\begin{bmatrix}7\\ 6\end{bmatrix}$$\begin{bmatrix}3\\ 4\end{bmatrix}$$\begin{bmatrix}7\\ 6\end{bmatrix}$
Figure 1: Example of a compositional reasoning question. Inside each node is the vector corresponding to this node. Red lines indicate the answer of each sub-question, given by $\arg\max_{u\in\mathcal{C}(v)}\{\mathbf{x}_{v},\mathbf{x}_{u}\}$, where $\mathcal{C}(v)$ are the children of node $v$.
2 Related Work
Expressive power of transformers.
[…]
Past work has also studied the expressive power of transformers with logarithmic-depth. (Sanford et al., 2024a, b) focus on graph tasks, showing how the parallelism of transformers gives them an advantage over other architectures. Liu et al. (2022) show the ability of logarithmic-depth transformers to simulate finite automata, and Merrill and Sabharwal extend these results by weakening their assumptions. We note that as implied by our Lemma 4.2, CRQs in general cannot be solved by finite-state automata.
Power of RNNs
An older line of research studied the capacities of recurrent networks, usually by relating them to automata (Merrill, 2019; Korsky and Berwick, 2019; Merrill et al., 2020). However, they generally assume that the order of the inputs is fixed and that the size of the network is constant in the size of the input. We study models whose hyperparameters (size, number of CoT tokens) change with the problem size, and we study the effect of both benevolent and adversarial orderings of the inputs.
Chain of Thought prompting.
Chain of thought (CoT) (Reynolds and McDonell, 2021; Wei et al., 2022b; Nye et al., 2021) is a method that enhances the ability of transformers to solve logical reasoning tasks. Rather than producing the answer all at once, the model is allowed to autoregressively generate a series of intermediate tokens to help it carry out each step in the solution. Li et al. (2024) show that constant-depth transformers that generate $T$ CoT tokens can simulate Boolean circuits of size $T$. Merrill and Sabharwal (2023b) prove that they can simulate Turing machines that run in $O(T)$ steps.
[…]
3 Problem Formulation and Preliminaries
Notations.
We use bold-face letter to denote vectors, e.g. $\mathbf{x},\mathbf{y}$. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and $1\leq i<j\leq d$ we define by $(\mathbf{x})_{i:j}\in\mathbb{R}^{j-i+1}$ the vector whose entries are the entries of $\mathbf{x}$ in places $i$ until $j$. We also define $(\mathbf{x})_{-1}$ the last coordinate of $\mathbf{x}$. For an integer $n$ we denote by $[n]=\{1,\dots,n\}$.
Let $T=(V,E)$ be a rooted tree with root $v_{r}\in V$. The depth of a node $v\in V$ is the length of the unique path from $v\in V$ to $v_{r}$. The depth of the tree is the largest depth of its nodes. The parent of a node $v\in V$ is the node connected to $v$ on the path to the root and denoted by $\mathcal{P}(v)$. The children of $v$ are the nodes whose parent is $v$, and are denoted by $\mathcal{C}(v)$. The degree of a node is the number of its children. Thus, leaves have a degree of 00. We say that a node $u$ is a descendant of $v$ if there is $i$ such that $\mathcal{P}^{(i)}(u)=v$, where $\mathcal{P}^{(i)}$ is the parent function composed with itself $i$ times. We also say in this case that $v$ is an ancestor of $u$.
3.1 Compositional Reasoning Questions
We formally define the problem of Compositional Reasoning Questions as follows:
Definition 3.1.
A Compositional Reasoning Question (CRQ) is a rooted tree $T=(V,E)$ with root $v_{r}\in V$, where each node in the tree $v\in V$ is labeled by a vector $\mathbf{x}_{v}\in\Gamma^{d}$. Here $\Gamma\subset\mathbb{R}$ is some finite vocabulary of constant size and $d\in\mathbb{N}$.
The size of the CRQ is defined as $|V|$. The answer to the CRQ is defined as $\mathcal{A}(v_{r})$ where the function $\mathcal{A}$ is defined recursively: For a leaf $u\in V$ we define $\mathcal{A}(u)=\mathbf{x}_{u}$. For a non-leaf node $v\in V$ we define:
$\mathcal{A}(v)=\operatorname*{arg\,max}_{\mathbf{x}_{u},u\in\mathcal{C}(v)}\left\langle\mathcal{A}(u),\mathbf{x}_{v}\right\rangle~{}.$
We refer to every node $v\in V$ which is not a leaf as a sub-question.
In the above definition we assume that the CRQs are defined over some finite vocabulary $\Gamma$. We can think about the vocabulary as $\Gamma=\{0,\pm 1,\dots,\pm 9\}$ for simplicity, although other vocabularies can be used (e.g. $\Gamma=\{\pm 1\})$.
CRQs are thus inherently hierarchical tasks, where the usual sequential structure is replaced by a tree structure. As a result, we will consider learning CRQs with models that can leverage this tree structure. Each node $\mathbf{v}$ is embedded as the vector $(\mathbf{x}_{v},\mathbf{z}_{v},\mathbf{z}_{\mathcal{P}(v)},\ell(v))$, where $\mathbf{x}_{v}$ is the label, $\mathbf{z}_{v}$ is a positional embedding, and $\ell(v)$ is the depth of $v$ in the tree.
3.2 Transformers
We now provide a formal definition of the transformer architecture that will be used throughout the paper. The input sequence is $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in\mathbb{R}^{d}$. We concatenate Positional Encodings (PE) to the inputs, which have the form $\mathbf{e}_{1},\dots,\mathbf{e}_{n}\in\mathbb{R}^{e}$. The positional encodings cannot depend on the values of the vectors $\mathbf{x}_{i}$, only on their positions (namely, on $i$). The input to the transformer is the sequence $\tilde{\mathbf{x}}_{i}=\begin{pmatrix}\mathbf{x}_{i}\\ \mathbf{e}_{i}\end{pmatrix}\in\mathbb{R}^{d+e}$.
The transformer backbone contains alternating layers of hardmax, single-head self-attention and feed forward networks with skip connections. For simplicity, we do not use layer normalization or attention masking (though the lower bounds generally hold against models that include them). Formally, let $\bm{h}_{i}^{(\ell)}$ denote the hidden embedding output after $\ell$ transformer blocks corresponding to the $i$th input. Define:
$\bm{h}_{i}^{(0)}$ $=\tilde{\mathbf{x}}_{i}$ (1)
$\bm{h}_{i}^{(\ell+1/2)}$ $=\bm{h}_{i}^{(\ell)}+\bm{V}_{\ell}\operatorname*{arg\,max}_{\bm{h}\in\{\bm{h}_{1}^{(\ell)},\ldots,\bm{h}_{n}^{(\ell)}\}}\left({\bm{h}}^{\top}\bm{K}_{\ell}^{\top}\bm{Q}_{\ell}\bm{h}_{i}^{(\ell)}\right)$ (2)
$\bm{h}_{i}^{(\ell+1)}$ $=\mathrm{MLP}_{\theta_{\ell}}\left(\bm{h}_{i}^{(\ell+1/2)}\right)$ (3)
where $\bm{V}_{\ell},\bm{K}_{\ell},\bm{Q}_{\ell}\in\mathbb{R}^{(d+e)\times(d+e)}$ and $\theta_{\ell}$ are weight matrices and $\sigma(\cdot)$ is the ReLU function. Ties in the $\operatorname*{arg\,max}$ operation are broken arbitrarily.
Finally, an unembedding layer discards entries corresponding to the positional encoding:
$\mathbf{y}_{i}=\begin{pmatrix}\bm{I}_{d}&\cdot\\ \cdot&\cdot\end{pmatrix}\bm{h}_{i}^{(L)}$ (4)
We consider $\mathbf{y}_{n}$ to be the output of the model. Following previous work (Merrill and Sabharwal, 2023a, c), we assume that the inputs, weights, and intermediate representations of the network can all be represented using $O(\log n)$ bits of precision.
4 Depth in Transformers is Necessary and Sufficient
In this section we provide two complementary results about the power of transformers to solve CRQs. We first show that deep transformers can solve any CRQ so long as the depth of the transformer is at least the depth of the CRQ tree. We then prove a conditional lower bound showing that constant depth transformers cannot solve all CRQs. Combining both results, we conclude that depth is both necessary and sufficient for transformers to solve CRQs.
4.1 Deep transformers can solve CRQs
This subsection presents our construction of a deep transformer that solves CRQs:
Theorem 4.1.
For any $L,n\in\mathbb{N}$ there exists a transformer $T$ with depth $L-1$ and embedding dimension $O(d+\log(n))$ that solves all the compositional reasoning questions up to $n$ nodes and depth $L$.
The proof appears in Section A.1. First, note that the depth of the transformer depends only on the depth of the tree of the CRQ. This is one of the main strengths of transformers: parallelism. Namely, the transformer is able to solve all the sub-questions in the tree of the same depth with just one layer.
We also emphasize that it is probably not possible to solve CRQs of size $n$ with less than $\log(n)$ precision. The reason is that a transformer without positional encoding is invariant to the order of the tokens. However, the CRQ has an inherent structure, a tree, which without it it cannot be solved. Thus, just to be able to provide even the simplest positional encoding which is numbering the nodes of the tree from $1$ until $n$ requires $O(\log(n))$ bits (cf. Merrill and Sabharwal (2023a)). Our construction uses a more complex positional encoding that captures the tree structure while still using only $O(\log(n))$ bits.
We now give a short intuition for our construction. We first define a positional encoding vector for each node that has four parts. For each non-leaf node $v$, we define an identifying vector $\mathbf{z}_{v}\in\{\pm 1\}^{O(\log(n))}$ such that $\left\langle\mathbf{z}_{v},\mathbf{z}_{u}\right\rangle\ll\left\lVert\mathbf{z}_{v}\right\rVert^{2}$ for any nodes $u,v\in V$. For leaf nodes, we let $\mathbf{z}_{v}=\boldsymbol{0}_{O(\log(n))}$. This is the first part. The second part of node $v$’s positional encoding vector is the identifier of its parent, $\mathbf{z}_{\mathcal{P}(v)}$. The third part is the depth of $v$ in the tree. The fourth is an indicator variable that is $1$ when the depth is $L-1$, where $L$ is the depth of the tree, and 00 otherwise. These four parts are all concatenated to the value of the node $\mathbf{x}_{v}$.
We construct each layer of the transformer to be exactly the same. Using the positional encoding, we construct the following attention pattern:
  • Each non-leaf tokens that have depth smaller than $L-1$ attends only to itself.
  • Each non-leaf tokens that has depth $L-1$ attends only to its children. Specifically, it attends to the child child that answers its sub-question, one whose value vector has the largest correlation with the non-leaf token.
  • Leaf tokens can do as they like, as they will not be attended to in succeeding levels.
By constructing this attention pattern, all the sub-questions in layer $L-1$ are solved simultaneously. We then use the MLP to reorganize the tokens. We increase the counter of the depth of every token, and update the indicator for nodes of depth $L-1$. Nodes that in the previous iteration had depth $L-1$ are altered to resemble leaves, i.e. their $\mathbf{z}_{v}$ vector is set to zero. For these nodes, we also set their value vector $\mathbf{x}_{i}$ to be the answer to their sub-question, as computing in the previous attention layer. In the next attention layer, nodes previous in layer $L-1$ will play the part of leaves and be attended to by their parents, which were previously in layer $L-2$. Applying this construction $L-1$ times will compute the answer to the root, which is the answer to the CRQ.
4.2 Constant depth transformers cannot solve CRQs
[…]
Lemma 4.2.
[…]
Theorem 4.3.
[…]
We note that arithmetic problems, as presented in Feng et al. (2024), can also be embedded as certain generalized CRQs. This embedding is presented in Section 7 and Appendix F. Thus, both Theorem 4.1 and Theorem 4.3 can be applied to arithmetic problems, if they are presented in a tree structure. It is in general possible to apply a preprocessing procedure using a depth $O(\log(n))$ transformer to turn an arithmetic problem presented in a sequential form, into a tree structure which captures the order of arithmetic operations.
5 Solution Using Shallow RNNs
In this section we will show that under certain assumptions, RNNs can solve CRQs. The proofs for this section appear in Appendix C. We first define RNNs in the following way:
Definition 5.1.
An RNN is a fully-connected neural network $\mathcal{N}:\mathbb{R}^{d+m}\rightarrow\mathbb{R}^{m}$. The inputs to the RNN are $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in\mathbb{R}^{d}$, and the RNN operates as $\mathcal{N}\left(\begin{pmatrix}\mathbf{x}_{i}\\ \mathbf{h}_{i-1}\end{pmatrix}\right)=\mathbf{h}_{i}$ where $\mathbf{h}_{0},\dots,\mathbf{h}_{n}\in\mathbb{R}^{m}$. The $\mathbf{h}_{i}$’s are called the hidden states of the RNN. $\mathbf{h}_{0}$ is defined as part of the architecture, and independent of the input data. The output of the RNN is $\mathbf{h}_{n}$ [^1] It is often common to define two separate outputs of the RNN, the hidden state and the prediction. Here we combine them for simplicity..
Note that in contrast to transformers, RNNs are not invariant to changing the order of the inputs. The reason is that the order in which the vectors $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ are fed to the model is part of the input itself, and not part of the architecture like positional encodings for transformers. We will show that the way the nodes of the CRQs are ordered is crucial for having an efficient solution with RNNs.
5.1 The order of the nodes matters
Our next result shows that there are bad orderings. This means that we can construct a family of CRQs and order the nodes adversarially so that any RNN must have a hidden dimension of size $\Omega(n)$ to solve them successfully.
Theorem 5.2.
Let $T=(V,E)$ be a balanced binary tree of size $n$. There exists an ordering of the nodes with the following property: Any RNN that reads the inputs in that order and solves all CRQs defined on $T$ must have a hidden state of size $\Omega(n)$ bits. That is, if the hidden state is in $\mathbb{R}^{m}$ and each entry is represented by $p$ bits, then $p\cdot m=\Omega(n)$.
The proof is given in Section C.1. Note that Theorem 5.2 doesn’t depend on the depth of the RNN. This means that given certain ordering of the nodes, the RNN must encode many of them into the hidden dimension. Also note that it is easy to construct an RNN with a hidden dimension of size $O(n)$ that solves all CRQs. The idea is just to encode all the nodes in the hidden state, which can be done if we have enough memory, and then use a large enough neural network to solve the CRQ. Finding such a neural network is also doable because $2$-layer networks have the universal approximation property (Cybenko, 1989; Leshno et al., 1993). Since the number of possible outputs is finite (because the alphabet is finite), we can approximate the solution up to a small enough accuracy and then threshold over the output.
The proof of Theorem 5.2 (Section C.1) uses a communication complexity argument. Specifically, we use a reduction from the set disjointness problem (see C.1 for a formal definition). We then construct a communication protocol between two parties, where each one of them has knowledge of only half of the inputs. When using the bad ordering of the nodes, this communication protocol forces the first party to encode all of its inputs into the hidden state when passing it to the second party.
We also note that Feng et al. (2024) proved that a specific construction of RNNs that solve arithmetic problems requires a hidden state of size $\Omega(n)$. Our results are stronger in the sense that: (1) Our lower bound is general and applies to any possible construction, so long as the bad order is used, and (2) We next show that the $\Omega(n)$ memory requirement can be greatly reduced just by re-ordering the inputs.
5.2 Memory-rank sort
In this subsection, we restrict ourselves to CRQs with full binary trees. That is, the degree of each node is 00 or $2$. This simplifies the proofs and presentation. In Appendix D, we explain how to extend the results from this section to non-binary trees. Also, note that every CRQ with a max degree of $k$ for each node can be converted into an equivalent CRQ with a binary tree, and the number of nodes increases by a factor of at most $2$ (see Section D.1). We also note that a node of degree $1$ is unnecessary in the context of CRQs, because a sub-question of degree $1$ has only one possible answer.
We will now introduce a sorting algorithm for trees that will allow us to solve any CRQs using an RNN with a small hidden dimension. First, we to define the memory-rank of each node:
Definition 5.3 (Memory Rank).
Let $T=(V,E)$ be a rooted tree. The memory rank of a node $v\in V$ is defined recursively as: $\textbf{mr}(v)=0$ if $v$ is a leaf, and otherwise
$\textbf{mr}(v)$ $=\max\left(a_{\max},a_{\min}+1\right)$
$\mathrm{where}\quad a_{\max}$ $=\max\left(\textbf{mr}(c_{1}(v)),\textbf{mr}(c_{2}(v))\right)$
$a_{\min}$ $=\min\left(\textbf{mr}(c_{1}(v)),\textbf{mr}(c_{2}(v))\right)$
and $c_{1}(v)$ and $c_{2}(v)$ are the two children nodes of $v$. The memory rank of the tree $\textbf{mr}(T)$ is defined as the memory rank of its root.
Intuitively, the memory rank is the smallest possible stack size needed for a stack machine to solve the CRQ if we are allowed to pick the order of the nodes. A stack machine is an automaton augmented with a stack that it can uses as memory. It processes the inputs sequentially, optionally pushing and popping a finite number of elements from its stack at each step and performing computations with the resulting vectors. We should order the inputs according to a post-ordering depth-first search of the CRQ, meaning that the position of a node in the ordering corresponds to when the depth first search last visits it. This ordering akin to reverse Polish notation, which notates an arithmetic expression like $(2+3)*6$ as the following sequence: 2 3+ 62\,3\,+\,6\,*2 3 + 6 ∗. By ordering the inputs this way, we ensure that whenever the automaton reads a non-leaf node, the top two elements on the stack are the solutions to the CRQs defined by that node’s two children. It can simply pop them off, compute the solution to the current node, and push the result back onto the stack to save it for later.
Even using a depth-first search ordering, the size of the stack used by this machine could grow as large as $n$ in the worst case. However, by ordering the children of each node carefully, we can reduce the stack depth to only $\log n$. To compute the CRQ corresponding to a given node $v$, we (1) solve for one of its children, (2) save this result on the stack, then (3) solve for the second child. If we solve $c_{1}(v)$ first, then the largest stack depth during the first stage is $\textbf{mr}(c_{1}(v))$ and the largest stack depth during the second stage is $\textbf{mr}(c_{2}(v))+1$. The overall largest stack depth is the maximum of these two. Thus, to use as little memory as possible, we should always start with the child whose memory rank is larger. This sorting algorithm is defined formally in Algorithm 1. See Figure 2 for an example of memory-rank sort of a tree. In Theorem 5.4, we will simulate this stack machine using a RNN with hidden dimension of size $\textbf{mr}(T)$.
211010002100100151471213361110124589
Figure 2: Memory-rank sorting of a tree. Each node is labeled with its memory rank, which ranges from 0 to 2 in this case. The number below each node is its ordering according to Algorithm 1. Note that we first traverse the right branch of the tree, since the right child of the root has a higher memory-rank than the left child.
5.3 Shallow RNNs can solve CRQs
We are now ready to present the main theorem of this section:
Theorem 5.4.
For any $n\in\mathbb{N}$ there exists an RNN with $5$ layers and hidden dimension $O(d\log(n))$ that solves any CRQ with a binary tree, if the nodes are ordered by Algorithm 1
The proof is given in Section C.2. The main idea of the proof is to simulate a stack machine with the RNN. The stack will contain the vectors $\mathbf{x}_{i}$ that are needed for future calculations, and will have two possible operations: (1) pop out vectors, and (2) push a vector. The stack will be transferred through the RNN in the hidden state. We include the depth of each node in its positional embedding, which will help us determine which of the two operations (push or pop) to use.
At each new input we check whether the last node in the stack has a larger depth by exactly $1$. If it doesn’t, then we insert $v$ into the stack and move to the next input. If it does, then we pop out the last two nodes from the stack, say $u_{1}$ and $u_{2}$, and calculate the inner products $\left\langle\mathbf{x}_{v},\mathbf{x}_{u_{1}}\right\rangle,\left\langle\mathbf{x}_{v},\mathbf{x}_{u_{2}}\right\rangle$. We insert back to the stack either $\mathbf{x}_{u_{1}}$ or $\mathbf{x}_{u_{2}}$, whichever has a higher inner product, but with the depth of the node $v$. All the above operations can be simulated using ReLU neural networks. The memory-rank sort does the rest of the work, since it makes sure that if $v$ is the current node to be processed, the last two nodes in the stack must either be its children, or nodes with a smaller depth. Doing this recursively over all the nodes will gradually solve the CRQ.
[…]
Finally, note that the memory rank of any tree with $n$ nodes is bounded by $\log(n)$ (see Lemma C.2). In fact, for a given tree structure $T$, the hidden dimension of the RNN needed to solve all CRQs with this structure is bounded by $O(d\cdot\textbf{mr}(T))$ using our construction. We next prove that the memory-rank also provides a lower bound on the required size of the hidden state of the RNN to solve CRQs:
Theorem 5.5.
Let $n\in\mathbb{N}$, let $T$ be a rooted tree of size $n$ with some ordering of the nodes from $1$ to $n$. Suppose there exists an RNN that solves all CRQs with a tree structure of $T$ if the nodes are provided in the given ordering. Then, the hidden dimension of the RNN must have $\Omega(\textbf{mr}(T))$ bits. In particular, an RNN that solves all CRQs on all trees of size $n$ for a given ordering must have a hidden dimension with $\Omega(\log(n))$ bits.
The proof is given in Section C.3. Combining Theorem 5.4 and Theorem 5.5, we see that the memory-rank provides a complete characterization of the memory required to solve CRQs using RNNs.
6 Solution Using Shallow Transformers with Chain of Thought
In this section we show that adding the ability to produce chain of thought tokens can also help solve CRQs using a constant depth transformers. We assume in this section that the nodes are ordered in reverse BFS ordering. Namely, for a tree with $n$ nodes and depth $L$, the nodes are numbered from $1$ until $n$ starting from the nodes at depth $L$, and going up until the root which is numbered $n$. Our main result in this section is the following:
Theorem 6.1.
There exists a $2$-layer transformer that solves all the CRQs with trees containing $n$ nodes. The embedding dimension is $O(d+\log(n))$, the bit-precision of the transformer is $O(\log(n))$, and the number of chain-of-thought tokens generated is $n$.
The proof idea is to solve all the sub-questions of the CRQ from the bottom-up. Each generated CoT token will be a solution to one of the sub-questions represented by a non-leaf node. Each node token contains two sets of positional encoding. The first one represents the tree structure and is similar to the positional encoding from Theorem 4.1. The second one encodes the position in the reverse BFS ordering of the nodes. The role of this embedding is so that at each operation of the transformer, the last token that was generated will attend to the next token in the reverse BFS order. This way, all the sub-question in the same layer are solved one after the other, before moving on to the next layer. The construction of the transformer itself contains $2$-layers of self-attention. The first layer is similar to a single layer of the construction from Theorem 4.1. Its role is to solve a single layer of sub-questions. The second layer make sure that the last generated token can only attend to the next token in the reverse BFS order, using the designated positional encodings.
[…]
We now compare our result with the CoT solution from Feng et al. (2024) for arithmetic problems. Both solutions use a transformer with $O(1)$ layers and $O(\log(n))$ bit precision. However, their solution generated $O(n^{2})$ [^2] In the arithmetic problems from Feng et al. (2024) we refer to $n$ as the number of symbols in the problems. i.e. numbers, parenthesis and arithmetic operations. CoT tokens, while ours generates only $O(n)$ tokens. This is a significant improvement, since for large sequences, a quadratic number of generated tokens can be infeasible to generate during inference time. The reason for this improvement is that in our solution each generated token solves exactly one sub-question, thus the number of tokens can be bounded by the number of sub-questions. In the solution of Feng et al. (2024), each time one sub-problem is solved (e.g. adding two numbers) the entire expression is rewritten. Thus, the number of tokens needed is $n+(n-1)+(n-2)+\cdots=O(n^{2})$. However, we use a different input format that encodes the tree structure directly into the positional encodings, rather than processing the arithmetic expression from left to right.
7 Generalization to Different Tasks
In our definition of CRQs (Definition 3.1) each sub-question is defined as the argmax over inner products. There are two main reasons for this choice: (1) The argmax function aligns well with the attention layers in transformers, allowing for simple constructions, and (2) CRQ is a flexible framework that includes natural problems like Boolean expressions as special cases. However, the CRQ framework can be modified to include many more tasks, which we term general CRQ (Definition F.1). Our upper bounds can be adapted to general CRQs. Namely, consider $k$ operators $f_{1},\dots,f_{k}:\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$, where each operator can be implemented by a ReLU network. We can now define CRQs over binary trees, where each non-leaf $v$ node is labeled by both a vector $\mathbf{x}_{v}$ and an operator $f_{i}$ for some $i\in[k]$, while leaf nodes are labeled only by a vector. The answer to the sub-question of a node $v$ with leaf children $u_{1},u_{2}$ will be $f_{i}(\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}})$.
To give a concrete example of the generality of this formalism, it allows to represent arithmetic problems. Let $d=1$ and consider the operators $f_{+},f_{-},f_{\times},f_{\div}$ defined as $f_{+}(\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}})=\mathbf{x}_{u_{1}}+\mathbf{x}_{u_{2}}$ (ignoring the vector $\mathbf{x}_{v}$) and analogously for the other operators. Now, an arbitrary arithmetic formula can be parsed into a tree and reformulated as a general CRQ. In Appendix F we provide a more extensive discussion about these generalizations, and explain how our constructions can be readily adapted to solve them.
8 Discussion and Future Work
In this work we present a general framework of problems which we term compositional reasoning questions. We show the depth trade-offs of solving CRQs using deep transformers, RNNs, and shallow transformers with CoT. Our results indicate that although transformers are highly parallelizable, they must be deep being able to solve CRQs with depth scaling with the depth of the tree of CRQs, this depth requirement is necessary. On the other hand, RNNs and CoT can solve CRQs with constant depth, although their solution is not parallelizable, and requires sequential computation.
There are several open questions remaining. First, is our solution using constant depth transformers that generates $n$ CoT tokens optimal, or is it possible to solve CRQs while generating fewer tokens? Second, our work only considers the expressiveness point of view, it would be interesting to understand the optimization process of transformers and other models when solving CRQs. Finally, it would be interesting to understand how general the solutions learned by transformers are. Do they generalize to other trees with different structures, sizes and different question regimes?
Acknowledgments
This work was partially supported by the Alfred P. Sloan Foundation, awards NSF RI-1816753, NSF CAREER CIF 1845360, NSF CHS-1901091, NSF DMS-MoDL 2134216, and an NSF Graduate Research Fellowship. We thank Will Merrill for useful discussions while this work was being completed.
Appendix A Proofs from Section 4
A.1 Proof of Theorem 4.1
The following lemma is useful for defining the positional encodings:
Lemma A.1.
For any $k\geq 2$ there exist $\mathbf{v}_{1},\dots,\mathbf{v}_{k}\in\{\pm 1\}^{4\log(k)}$ such that $|\left\langle\mathbf{v}_{i},\mathbf{v}_{j}\right\rangle|\leq 3\log(k)$ for all $i\neq j$.
Proof.
We sample $\mathbf{v}_{1},\dots,\mathbf{v}_{k}\in\{\pm 1\}^{4\log(k)}$ uniformly at random. Note that for any $i\neq j$ we have that $\mathbb{E}[\left\langle\mathbf{v}_{i},\mathbf{v}_{j}\right\rangle]=0$. By Hoeffding’s inequality we have that:
$\text{Pr}\left(\left|\left\langle\mathbf{v}_{i},\mathbf{v}_{j}\right\rangle\right|\geq 3\log(k)\right)\leq 2\exp(-4\log(k))~{}.$ (5)
Applying union bound on the above for all pairs $i\neq j$ we get that:
$\text{Pr}\left(\forall i\neq j,~{}\left|\left\langle\mathbf{v}_{i},\mathbf{v}_{j}\right\rangle\right|\leq 3\log(k)\right)\geq 1-2\exp(-4\log(k))\cdot k^{2}\geq 1-\frac{2}{k^{2}}~{}.$
In particular, for $k\geq 2$ this probability is non-zero, meaning that there exists such vectors $\mathbf{v}_{1},\dots,\mathbf{v}_{k}$. ∎
Theorem 4.1.
For any CRQ with up to $n$ nodes, there exist at most $|Y|\leq n$ possible queries. By Lemma A.1 there exist vectors $\mathbf{z}_{1},\dots,\mathbf{z}_{n}\in\{\pm 1\}^{4\log(n)}$ such that $|\left\langle\mathbf{z}_{i},\mathbf{z}_{j}\right\rangle|\leq 3\log(n)$ for any $i\neq j$.
We will first describe how we embed each node as an input token to the transformer, and then describe the construction of the transformer itself. For each node $v\in V$ that is not a leaf, we set one of the vectors $\mathbf{z}_{i}$’s to be corresponded with this node, and denote it as $\mathbf{z}_{v}$. Recall that $\mathcal{P}(v)$ is the parent of node $v$. Note that although each node may have several children, it only has one parent.
Each node of the tree will be embedded as a vector in $\mathbb{R}^{2d+8\log(n)+2}$. The input tokens corresponding to a leaf node $u$ and a non-leaf node $v$ are defined as:
$\mathbf{t}_{u}:=\begin{pmatrix}\mathbf{x}_{u}\\ \bm{0}_{d}\\ \mathbf{z}_{\mathcal{P}(u)}\\ \bm{0}_{4\log(n)}\\ 0\\ 0\end{pmatrix},~{}~{}\mathbf{t}_{v}:=\begin{pmatrix}\mathbf{x}_{v}\\ \bm{0}_{d}\\ \mathbf{z}_{v}\\ \mathbf{z}_{\mathcal{P}(v)}\\ \ell(v)\\ \mathds{1}(\ell(v)=L-1)\end{pmatrix}~{},$ (6)
where $\ell(v)$ is the depth of node $v$ (the length of the shortest path from $v$ to the root).
The crux of the construction lies mainly in the way we choose the embedding: The embedding first contains the value of the leaves or query of the corresponding node. The embedding also contains a positional encoding (PE), where each leaf token contains its parent’s PE, and each non-leaf token contains its own PE as well as its parent’s. In addition, there is a number representing the depth of each node and a flag on whether a node is in the second to last layer (i.e it is the deepest intermediate node connected to a leaf).
We construct the weights of the transformer so it will work as follows: In the first layer, each token representing an intermediate node will attend either to itself, if its depth is higher than $L-1$, or to its children, if its depth is exactly $L-1$. This means that only the tokens of nodes in depth $L-1$ will change while all other tokens of intermediate nodes will remain the same. Using the positional encodings, the tokens at depth $L-1$ will attend only to their children, not to themselves, and after the self-attention layer (before applying the $V$ matrix) will be equal to the token representing their child with the highest correlation to their query. We now use the $V$ matrix, residual connection and MLP so that the tokens with depth $L-1$ will have a similar embedding to the leaves tokens, and decrease the depth of all other intermediate tokens. We also make sure that the flag after the first layer will be equal to $1$ for nodes of depth $L-2$. We now apply a similar layer exactly $L$ times, so that in the final layer the output of the token representing the root will be equal to the solution to the CRQ.
We now turn to the formal construction of the transformer. The order of the input tokens is not important for the construction, except for the root token which will be the last token since its final embedding will include the answer to the CRQ. Let $c\geq 1$ be some universal constant such that $\left\lVert\mathbf{x}\right\rVert^{2}\leq c$ for every $\mathbf{x}\in\Gamma^{d}$, there exists such a constant since $\Gamma$ is finite. The matrices of all the layers of the transformer will be the same and equal to:
$K=\begin{pmatrix}I_{2d+8\log(n)}&&&\\ &0&&\\ &&1&\end{pmatrix},$
$Q=\begin{pmatrix}I_{2d}&&&&&\\ &6cI_{4\log(n)}&&&&\\ &&cI_{4\log(n)}&&&\\ &&&0&&&\\ &&&&-6c\log(n)&\end{pmatrix},$
$V=\begin{pmatrix}0_{d}&I_{d}&\\ I_{d}&0_{d}&\\ &&0_{8\log(n)+2}\end{pmatrix}~{}.$
Before defining the MLP, we will explain how the attention layer operates on the input tokens. Since the attention mechanism contains a hardmax head, the output of each token will depend only on the token to which it attends. We will first show the next two claims:
  1. 1.
    Each token non-leaf node $v$ with $(\mathbf{t}_{v})_{-1}=0$ will attend to itself.
  2. 2.
    Each token non-leaf $v$ with $(\mathbf{t}_{v})_{-1}=1$ will attend to $\mathbf{t}_{u}$ with $\arg\max_{\mathbf{t}_{u},u\in\mathcal{C}(v)}\left\langle\mathbf{x}_{u},\mathbf{x}_{v}\right\rangle$.
For the first claim, let $v$ be some node with $(\mathbf{t}_{v})_{-1}=0$. Then we have that:
$\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{v}=\left\lVert\mathbf{x}_{v}\right\rVert^{2}+6c\left\lVert\mathbf{z}_{v}\right\rVert^{2}+c\left\lVert\mathbf{z}_{\mathcal{P}(v)}\right\rVert^{2}\geq 28c\log(n)$ (7)
Now, let $u\neq v$ be some other node. If $u$ is a leaf then we have that:
$\mathbf{t}_{u}^{\top}K^{\top}Q\mathbf{t}_{v}$ $\leq|\left\langle\mathbf{x}_{u},\mathbf{x}_{v}\right\rangle|+6c|\left\langle\mathbf{z}_{\mathcal{P}(u)},\mathbf{z}_{v}\right\rangle|+c|\left\langle\boldsymbol{0},\mathbf{z}_{\mathcal{P}(v)}\right\rangle|$
$\leq c+24c\log(n)+<25c\log(n)~{}.$
If $u$ is not a leaf then we have that:
$\mathbf{t}_{u}^{\top}K^{\top}Q\mathbf{t}_{v}$ $\leq|\left\langle\mathbf{x}_{u},\mathbf{x}_{v}\right\rangle|+6c|\left\langle\mathbf{z}_{u},\mathbf{z}_{v}\right\rangle|+c|\left\langle\mathbf{z}_{\mathcal{P}(u)},\mathbf{z}_{\mathcal{P}(v)}\right\rangle|$
$\leq c+18c\log(n)+4c\log(n)<23c\log(n)~{},$
where we used that $|\left\langle\mathbf{z}_{u},\mathbf{z}_{v}\right\rangle|\leq 3\log(n)$ for $u\neq v$. This shows that $\mathbf{t}_{v}$ can only attend to itself.
Next, let $v$ be a node with $(\mathbf{t}_{v})_{-1}=1$. Let $u$ be a node with $u\in\mathcal{C}(v)$ and let $w\neq v$ with $w\notin\mathcal{C}(v)$. We have that:
$\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{u}$ $=\left\langle\mathbf{x}_{u},\mathbf{x}_{v}\right\rangle+6c\left\lVert\mathbf{z}_{v}\right\rVert^{2}$
$\geq-c+24c\log(n)\geq 23c\log(n)$
$\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{w}$ $\leq|\left\langle\mathbf{x}_{w},\mathbf{x}_{v}\right\rangle|+6|\left\langle\mathbf{z}^{\prime},\mathbf{z}_{v}\right\rangle|+|\left\langle\mathbf{z}_{\mathcal{P}(w)},\mathbf{z}_{\mathcal{P}(v)}\right\rangle|$
$\leq c+18c\log(n)+4c\log(n)<23c\log(n)$
$\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{v}$ $=\left\lVert\mathbf{x}_{v}\right\rVert^{2}+6c\left\lVert\mathbf{z}_{v}\right\rVert^{2}+c\left\lVert\mathbf{z}_{\mathcal{P}(v)}\right\rVert^{2}-6c\log(n)$
$\leq c+24c\log(n)+4c\log(n)-6c\log(n)<23c\log(n)~{}.$
Here $\mathbf{z}^{\prime}$ is some vector with $\mathbf{z}^{\prime}\neq\mathbf{z}_{v}$ (n fact, $\mathbf{z}^{\prime}=\mathbf{z}_{w}$ or $\mathbf{z}^{\prime}=\mathbf{z}_{\mathcal{P}(w)}$, depending on whether $\mathbf{w}$ is a leaf). This shows that $\mathbf{t}_{v}$ can only attend to its children. It is also clear that $\mathbf{t}_{v}$ will attend to its child $u$ that maximized $\left\langle\mathbf{x}_{u},\mathbf{x}_{v}\right\rangle$. This finishes the two claims above.
After applying the $V$ matrix and the residual connection from the previous layer, the output of the self-attention layer on token $\mathbf{t}_{v}$ is equal $\mathbf{t}_{v}+\mathbf{o}_{v}$, where $(\mathbf{o}_{v})_{(d+1:2d)}=\mathbf{x}_{v}$ if $(\mathbf{t}_{v})_{-1}=0$ and $(\mathbf{o}_{v})_{(d+1:2d)}=\arg\max_{\mathbf{x}_{u},u\in\mathcal{C}(v)}\left\langle\mathbf{x}_{u},\mathbf{x}_{v}\right\rangle$ if $(\mathbf{t}_{v})_{-1}=1$. In words, the entries of each token in place $d+1$ until $2d$ after the self-attention layer is the solution to the sub-question for the deepest non-leaves nodes of the tree, and for other nodes it is the queries themselves.
We now define an MLP that will reorganize the tokens before applying the next layer of self-attention. The role of the MLP will be to perform the following operations on an input token $\mathbf{t}_{v}$:
  1. 1.
    If the entries in places $12d+4\log(n)+$ 1 to $2d+8\log(n)$ are zero, then $\mathbf{t}_{v}=\boldsymbol{0}$.
  2. 2.
    If the last entry of $\mathbf{t}_{v}$ is equal to $1$, then replace the entries in places $12d+$ 1 until $2d+4\log(n)$ (which used to hold $\mathbf{z}_{v}$) with the entries in places $12d+4\log(n)+$ 1 until $2d+8\log(n)$ (which used to hold $\mathbf{z}_{\mathcal{P}(v)}$). Then, replace the entries in places $12d+4\log(n)+$ 1 until $2d+8\log(n)$ with zeroes. This makes it a leaf embedding.
  3. 3.
    Replace between the entries in places $1$ until $d$ and $d+1$ until $2d$. Then replace the entries in places $d+1$ until $2d$ with zeroes.
  4. 4.
    Add $1$ to the entry in place $12d+8\log(n)+$ 1. Then, if this entry is equal to $L-1$, replace the last entry with 1, if this entry is equal to $L$ replace the last two entries with zeroes.
The first operation zero out any token that corresponded to a leaf node, as they won’t be necessary in later stages of the computation. The second operation turn tokens in the second to last layer to have the same embedding as leaf nodes, namely that they only contain the PE of their parent, and don’t have a PE of their own. The third operation update the value of $\mathbf{x}_{v}$, this only changes for nodes that were in the second to last layer. For any other non-leaf node, since it attended to itself both vectors are equal. The last operation updates the depth of each node. Now, each node that was in the second to last layer will be embedded similarly to a leaf, while nodes in the third to last layer will have $1$ in their last coordinate.
Each such operation can be implemented by a $2$-layer MLP with width bounded by $O(d+\log(n))$. For example, to implement the first operation we use the function:
$f(z)=\sigma(z+1)-\sigma(z-1/2)+\sigma(1-z)-\sigma(1/2-z)~{},$
that can be implemented by the first layer of an MLP. We concatenate the output of this function on the $12d+4\log(n)+$ 1-th coordinate of the input, to the input itself. Then, we use the output layer to apply the function $\tilde{\mathbf{x}}\mapsto\boldsymbol{0}_{k}\cdot(\tilde{\mathbf{x}})_{k+1}+\tilde{\mathbf{x}}\cdot(1-(\tilde{\mathbf{x}})_{k+1})$, where $\tilde{\mathbf{x}}$ is the input concatenated with the additional coordinate of the output of $f$, and $k=2d+8\log(n)+2$. use similar constructions to implement the rest of the operations.
After applying the MLP we have that each token is either equal to: (1) $\boldsymbol{0}$ if it was previously a leaf. (2) An embedding of a leaf if it was previously in the second to last layer; or (3) An embedding of a non-leaf node if it wasn’t a leaf or in the second to last layer. By applying the same construction $L$ times, we get that each node except the root is equal to $\boldsymbol{0}$, while the first $d$ coordinates of the root contain the answer to the CRQ, which is the final output of the transformer. This finishes the proof. ∎
A.2 Proof of Lemma 4.2
$\begin{pmatrix}-1\\ -2\end{pmatrix}$$\begin{pmatrix}2\\ 1\end{pmatrix}$$\begin{pmatrix}-2\\ 1\end{pmatrix}$$\varphi$$\begin{pmatrix}1\\ 1\end{pmatrix}$$\begin{pmatrix}1\\ 0\end{pmatrix}$$\begin{pmatrix}0\\ 1\end{pmatrix}$
$\mathbf{t}_{0}/\mathbf{t}_{1}$$\varphi_{1}$$\varphi_{2}$
Figure 3: Left: Construction for $\neg\varphi$. Right: Construction for $\varphi_{1}\land\varphi_{2}$ and $\varphi_{1}\lor\varphi_{2}$
[…]
  1. 1.
    00 and $1$ are Boolean formula.
  2. 2.
    If $\varphi$ is a Boolean formula, then ($(\neg\varphi)$ is a Boolean formula.
  3. 3.
    If $\varphi_{1},\varphi_{2}$ are Boolean formulas, then $(\varphi_{1}\lor\varphi_{2})$ and $(\varphi_{1}\land\varphi_{2})$ are Boolean formulas.
The goal in BFEP, is given a Boolean formula to evaluate whether it is true or false (i.e. outputs $1$ or 00). We will construct a translation function $f$ from Boolean formulas to CRQ, and specifically to a tree structure where each node is given a specific value. We will use vectors in $\mathbb{R}^{2}$, while in fact all the possible values of the vectors in the CRQ will be in $\{0,\pm 1,\pm 2\}$. We define the vectors $\mathbf{t}_{1}=\begin{pmatrix}1\\ 0\end{pmatrix}$ and $\mathbf{t}_{0}=\begin{pmatrix}0\\ 1\end{pmatrix}$. The translation function is defined recursively as follows:
  1. 1.
    For $f(0)$ and $f(1)$ we construct a node in the tree whose corresponding value is equal to $\mathbf{t}_{0}$ and $\mathbf{t}_{1}$ respectively.
  2. 2.
    Given two Boolean formulas $\varphi_{1},\varphi_{2}$ we define $f((\varphi_{1}\land\varphi_{2}))$ and $f((\varphi_{1}\lor\varphi_{2}))$ as the tree in Figure 3 (right), where for $\land$ we use $\mathbf{t}_{0}$ for the root, and for $\lor$ we use $\mathbf{t}_{1}$.
  3. 3.
    Given a Boolean formulas $\varphi$ we define $f((\neg\varphi))$ as the tree in Figure 3 (left).
[…]
$r_{i}=\sum_{j<i}\mathds{1}\big{[}\varphi_{j}=`(\text{'}\big{]}-\sum_{j<i}\mathds{1}\big{[}\varphi_{j}=`)\text{'}\big{]}~{}.$ ’ ] . (8)
[…]
Appendix B Memory-rank sort
In this section we formally define the memory-rank sort. It is presented in Algorithm 1. We next prove that performing memory-rank sort and calculting the memory-rank of all the nodes in a tree can be done efficiently:
Lemma B.1.
Let $T=(V,E)$ be a rooted binary tree, then calculating the memory-rank of each node and performing the memory-rank sort takes time $O(|V|)$.
Proof.
Calculating the memory rank of each node can be done using a reverse breadth-first search (BFS) of the tree, traversing from the leaves to the root. BFS can be done in time $O(|V|)$, and calculating the memory rank of each nodes requires only knowledge of the memory rank of its children. Thus, this can be done in a linear time by going over each node only once.
The time that Algorithm 1 runs is the number of iterations in the “while” loop. Note that each leaf in the tree is visited at most once, since if $v_{\text{cur}}$ is a leaf, it is inserted to the list and not visited again. For any non-leaf node, it is visited at most $3$ times during the loop. Once for each one of its children, and again when it is inserted into the list. In total the running time of the sorting algorithm is bounded by $3|V|$. ∎
Input: A Tree $T=(V,E)$ with root $v_{r}$ and memory rank calculated for each node.
$S=[]$
$v_{\text{cur}}\leftarrow v_{r}$
while $|S|\neq|V|$ do
       if $\mathcal{C}(v_{\text{cur}})=\varnothing\textbf{ or }\mathcal{C}(v_{\text{cur}})\subseteq S$ then
            $S\leftarrow v_{\text{cur}}$
             $v_{\text{cur}}\leftarrow\mathcal{P}(v_{\text{cur}})$
             Continue
      $v_{\text{cur}}\leftarrow\underset{v\in\mathcal{C}(v_{\text{cur}}),v\notin S}{\arg\max}\textbf{mr}(v)$
Return: $S$
Algorithm 1 Memory-Rank Sort
Appendix C Proofs from Section 5
C.1 Proof of Theorem 5.2
Proof.
Our proof uses the following claim from communication complexity:
Claim C.1 (Lower bound for set disjointness [Yao, 1979]).
Suppose Alice and Bob are given inputs $a,b\in\{0,1\}^{n}$, respectively. Their goal is to calculate $\max_{i}a_{i}b_{i}$ by alternately sending $1$-bit messages to one another over a sequence of communication rounds. Any deterministic protocol for computing $\max_{i}a_{i}b_{i}$ requires at least $n$ rounds of communication.
For simplicity of the proof, we use trees with $14n-$ 1 nodes instead of $n$, this will only effect the constant in the $\Omega$ notation. We construct the following CRQ: The tree $T=(V,E)$ is a balanced binary tree with $14n-$ 1 nodes, thus it has $2n$ leaves. All the values of the leaves are in {±$\{\pm 1\}$. The nodes in the second to last layer are all equal to $-1$, and the rest non-leaf nodes are equal to $1$. We number the leaves as $v_{1},\dots,v_{2n}$. Suppose we are given an instance of the set-disjointness problem, where Alice is given the leaves $v_{2i-1}$ and Bob the leaves $v_{2i}$ for $i\in[n]$. By the definition of CRQ, the answer to the second to last layer nodes with leaves $v_{2i-1}$ and $v_{2i}$ is $1$ if $v_{2i-1}=v_{2i}=1$ and $-1$ otherwise. In addition the answer to any other node will be $1$ if one of its children is $1$ and $-1$ otherwise. In total, the answer to the tree is $1$ if there exists some $i\in[n]$ with $v_{2i-1}=v_{2i}=1$ and $-1$ otherwise.
Consider the following ordering of the nodes for the RNN: The first $n$ nodes are $v_{2i-1}$ for $i\in[n]$, the next $n$ nodes are $v_{2i}$ for $i\in[n]$. The root is the last node in this ordering, and the rest of the nodes are given in some arbitrary order. Suppose there is an RNN known to both Alice and Bob that solves the CRQ problem above in the prescribed order of the nodes, and it has a hidden state $\mathbf{h}\in\mathbb{R}^{m}$ where each coordinate is represented by $p$ bits. We will define the following communication protocol to solve the set disjointness problem: Alice apply the RNN on her inputs. After each input $v_{2i-1}$ for $i\in[n]$ she passes the hidden state $\mathbf{h}_{i}$ to the next recurrence of the RNN. After $n$ such recurrences of the RNN, the hidden state $\mathbf{h}_{n}$ is passed to Bob, he inputs his inputs $v_{2j}$ for $j\in[n]$ in the prescribed order, each time passing a hidden state. He continue to run the RNN on the rest of the nodes, and output the answer to the CRQ. The number of bits transferred between Alice and Bob is $m\cdot p$, since the only communication between them is transferring $\mathbf{h}_{n}$ and Alice finishes processing her inputs. By C.1 we have that $mp=\Omega(n)$, which finishes the proof.
C.2 Proof of Theorem 5.4
We first need the following two lemmas:
Lemma C.2.
For a binary tree $T$ with $n$ nodes, we have that $\textbf{mr}(T)\leq\log(n)$.
Proof.
Let $r_{k}$ be the smallest number of descendants for a node with memory-rank $k$. It is easy to see that $r_{1}=2$, which happens when the node has only two children that are leaves. Let $v$ be a node with $\textbf{mr}(v)=k$, and let $u_{1},u_{2}$ be its children. It can be seen that $r_{k}\geq 2r_{k-1}$. This is because, if $\textbf{mr}(u_{1})\neq\textbf{mr}(u_{2})$, then one of them must be equal to $k$, assume w.l.o.g it is $u_{1}$. This means that either $u_{1}$ or one of its descendants have two children with a memory-rank of $k-1$. In particular, the number of descendants of $v$ is larger than $2r_{k-1}$. If $\textbf{mr}(u_{1})=\textbf{mr}(u_{2})$, then they are both equal to $k-1$, in which case the same conclusion follows. Applying the recurrence formula, we get that $r_{k}\geq 2^{k}$.
Let $k$ be the memory rank of a tree with $n$ nodes. We have $2^{k}\leq r_{k}\leq n$. Thus, $k\leq\lfloor\log(n)\rfloor$. Hence, for any tree with $n$ nodes, the largest possible memory-rank of the tree is bounded by $\log(n)$.
Lemma C.3.
Let $R,\epsilon>0$ and $d\in\mathbb{N}$. There exists a $2$-layer neural network $\mathcal{N}:\mathbb{R}^{2d}\rightarrow\mathbb{R}$ with width $O\left(\frac{d^{2}R}{\epsilon}\right)$ such that $\max_{\mathbf{x},\mathbf{y}\in[-R,R]^{d}}|\mathcal{N}(\mathbf{x},\mathbf{y})-\left\langle\mathbf{x},\mathbf{y}\right\rangle|\leq\epsilon$
Proof.
We can use Lemma 6 from Daniely [2017] to find a $2$-layer network $N:\mathbb{R}^{2}\rightarrow\mathbb{R}$ with width $O\left(\frac{R}{\epsilon}\right)$ such that $\max_{x,y\in[-R,R]^{d}}|N(x,y)-x\cdot y|\leq\epsilon$. Summing $d$ such networks over the coordinates of $\mathbf{x}$ and $\mathbf{y}$, and replacing $\epsilon$ with $\epsilon^{\prime}=\frac{\epsilon}{d}$ proves the lemma. ∎
Theorem 5.4.
The main crux of the proof is to simulate a stack using a ReLU network with constant depth. The maximal size of the stack for a tree $T=(V,E)$ will be $O(d\cdot\textbf{mr}(T))$. The RNN will execute the following pseudo-code:
  1. 1.
    Given an input node $v$, check the last node $u$ in the stack:
    1. (a)
      If $u$ has the same depth or a larger depth than $v$, then add $v$ to the stack.
    2. (b)
      If $u$ has a depth that is lower by exactly $1$ than $v$, then pop out the last two nodes from the stack $u$ and $w$. If $\left\langle u,v\right\rangle\geq\left\langle w,v\right\rangle$ then insert $u$ into the stack and raise its depth by $1$. Otherwise, inset $w$ and raise its depth by $1$.
  2. 2.
    The stack is the hidden state that is transferred to the next iteration of the RNN.
We now turn to the formal construction. Each node $v\in V$ in the tree will be embedded as a vector $\begin{pmatrix}\mathbf{x}_{v}\\ \ell(v)\end{pmatrix}\in\mathbb{R}^{d+1}$, where $\ell(v)$ is the depth of $v$. The hidden state of the RNN at the beginning will be $\mathbf{h}=\bm{0}_{d(\log(n)+1)}$. The input to the RNN will be the current input vector, concatenated to the hidden state. Throughout the proof we refer to $\tilde{\mathbf{x}}$ the vector that is inputted to the RNN, namely the concatenation of the embedding of each node with the hidden state. We call the part of $\tilde{\mathbf{x}}$ that contains the hidden state $\mathbf{h}$ the stack, and each consecutive $d+1$ coordinates in it as a node that is was inputted to the stack. The reason is that this part of the input will simulate a stack where the RNN can only access its first and second nodes, namely the first $22d+$ 2 coordinates. The RNN will add at most $3$ coordinates to $\tilde{\mathbf{x}}$, which will simulate flags with values that are either 00 or $1$. The RNN contains the following layers:
  1. 1.
    Layer 1: Check whether the depth of the input node $v$ is the same or larger than the depth of the last node in the stack. If so, turn on a specific flag for this event.
  2. 2.
    Layer 2: If the flag is turned on, insert $v$ into the stack by moving all the elements $\log(n)+1$ entries forward and putting $v$ at the top of the stack.
  3. 3.
    Layer 3: If the flag is turned off, extract the last two nodes from the stack $u_{1}$ and $u_{2}$, and calculate $\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle,~{}\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle$.
  4. 4.
    Layers 4 & 5: If the flag is turned off, insert $u_{1}$ or $u_{2}$ to the stack, whichever has the higher inner product with $v$. Also, increase the depth of this node by $1$.
Note that the input to the RNN is of dimension $d(\log(n)+1)+d+1$, which is a concatenation of the embedding of the current input node and the hidden state. We will also need to following parameters: Let $c>0$ be such that $c>|a|$ for every $a\in\Gamma$, namely, it is larger than all possible entries of the $\mathbf{x}_{v}$’s, and that $c>L$, the depth of the tree. We have that $c=O(n)$ which is the largest possible depth of the tree. Also, let $\epsilon>0$ such that $\min_{a,b\in\Gamma}|a-b|<\epsilon$. We think about $\epsilon$ as a constant independent of $n$ and $d$, as it only depends on the possible values in $\Gamma$. Each layer is constructed in the following way:
First layer: We denote the input vector to the RNN is $\tilde{\mathbf{x}}$. The first layer add a coordinate to the input. All the input coordinates are copied, while the last coordinate is constructed as: $\sigma((\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1}+1)-\sigma((\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1})$. Note that this added last coordinate is equal to $1$ if $(\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1}\geq 0$, meaning that the current node has an equal or larger depth that the last node in the stack, and 00 if $(\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1}\leq-1$. The copying of the input can be done by a ReLU network, since $\sigma(z)-\sigma(-z)=z$ for every $z\in\mathbb{R}$. Thus, adding two identity matrices to the weights of the MLP with different signs, and adding their outputs together will copy the inputs.
Second layer: Let $A=\begin{pmatrix}I_{d+1}&\boldsymbol{0}\\ I_{d(\log(n)+1)}&\boldsymbol{0}\end{pmatrix}$ be a $d(\log(n)+1)+d+1$ square matrix. The second layer will perform the following operation:
$\mathbf{x}\mapsto\sigma\left(I\mathbf{x}_{(1:-2)}-2cI\mathbf{x}_{(-1)}+c\boldsymbol{1}\right)+\sigma\left(A\mathbf{x}_{(1:-2)}-2cI(-\mathbf{x}_{(-1)}+1)+c\boldsymbol{1}\right)-c\boldsymbol{1}~{}.$
This operation can be implemented by a $2$-layer network, where the constant $\boldsymbol{1}$ vectors are added using the bias terms. This operation will apply the identity matrix to the inputs if the flag (defined in the previous layer) is 00, and apply the $A$ matrix if it is $1$. The $A$ matrix will copy the first $d+1$ entries (the current input node), and will move also add it to the stack, while also moving all the other entries $d+1$ coordinates further down in the stack. The added $c\boldsymbol{1}$ factor is so that if there are negative values in the entries of the inputs, they will not be removed by the ReLU, this factor is removed at the end to keep the original value. We will later on prove that the stack will not get overflown, meaning that we don’t delete the last $d+1$ entries of it.
Third layer: This layer will be used to approximate inner products. Assume that $v$ is a non-leaf node and let $u_{1}$ and $u_{2}$ be its two children. If it is a leaf node, this layer will not effect its input. The layer will add two coordinates to its input vector, corresponding to an approximation of $\left\langle\mathbf{x}_{u_{1}},\mathbf{x}\right\rangle$ and $\left\langle\mathbf{x}_{u_{2}},\mathbf{x}\right\rangle$.
By Lemma C.3 there exists a $2$-layer neural network $\mathcal{N}:\mathbb{R}^{2d}\rightarrow\mathbb{R}$ with width $O(d^{2})$ such that $\max_{\mathbf{x},\mathbf{y}\in\Gamma^{d}}|\mathcal{N}(\mathbf{x},\mathbf{y})-\left\langle\mathbf{x},\mathbf{y}\right\rangle|\leq\frac{\epsilon}{4}$. We will stack two such networks that approximate up to an error $\frac{\epsilon}{4}$ the inner products $\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle,\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle$. If $\mathbf{x}_{u_{1}}\neq\mathbf{x}_{u_{2}}$, then by the assumption on $\epsilon$ we have that $|\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle-\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle|\geq\epsilon$. By the construction of $\mathcal{N}$, we also have that:
$|\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle-\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle|\leq|\mathcal{N}(\mathbf{x}_{u_{1}},\mathbf{x}_{v})-\mathcal{N}(\mathbf{x}_{u_{2}},\mathbf{x}_{v})|+\frac{\epsilon}{2}~{}.$
This means that $|\mathcal{N}(\mathbf{x}_{u_{1}},\mathbf{x}_{v})-\mathcal{N}(\mathbf{x}_{u_{2}},\mathbf{x}_{v})|\geq\frac{\epsilon}{2}$. We use a similar construction to the previous layer to apply this calculation and add $\mathcal{N}(\mathbf{x}_{u_{1}},\mathbf{x}_{v})$ and $\mathcal{N}(\mathbf{x}_{u_{2}},\mathbf{x}_{v})$ only if the flag in the last coordinate is turned off. Otherwise, if the flag is turned on, we just copy the inputs.
Fourth and Fifth layers: In the last layer, we add an additional flag on whether $\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle>\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle$ and 00 otherwise. This can be calculated by using the following function:
$\mathbf{x}\mapsto\frac{2}{\epsilon}\left(\sigma((\mathbf{x})_{-2}-(\mathbf{x})_{-2})-\sigma((\mathbf{x})_{-2}-(\mathbf{x})_{-2}-\frac{\epsilon}{2})\right)~{}.$
By the argument from the previous layer, this function will provide the additional flag. All the other coordinates are copied as is. We use another layer to pop out the first $22d+$ 2 coordinates from the stack, meaning that all the entries in the stack are moved upward by $d+1$ coordinates, and the first $d+1$ are changed to either $\begin{pmatrix}\mathbf{x}_{u_{1}}\\ \ell(u_{1})-1\end{pmatrix}$ or $\begin{pmatrix}\mathbf{x}_{u_{2}}\\ \ell(u_{2})-1\end{pmatrix}$, depending on whether the new flag is $1$ or 00. Again, we apply this operation only if the flag from the first layer is turned off, by using a similar construction to the second layer. The rest of the coordinates are copied, and the coordinates of the flags that were used throughout the computation are removed. The output will be the hidden state to the next recurrence with the dimension as the hidden state that was inputted.
Proof of correctness: First, it is clear from the construction that applying the RNN on a single input node will perform the pseudo-code described in the beginning of the proof. We will show the the construction outputs the correct answer to the CRQ. When a node is inputted to the RNN, according to the ordering of the memory-rank sort, there are three options:
  1. 1.
    Its depth is equal to the the depth of the first node in the stack. This can only happen to leaf nodes, in which case, according to the memory-rank sorting, the node in the stack must be a sibling of the inputted node.
  2. 2.
    Its depth is larger than the depth of the first node in the stack. In this case, the first node in the stack will be pushed down to the second place after applying the RNN.
  3. 3.
    Its depth is smaller by exactly $1$ than the last node in the stack. In which case, it must be its parent, and the second node in the stack is its sibling. After applying the RNN, both nodes are popped out, and the node that correctly answer the query corresponding to $v$ is inputted back to the stack, but with an updated depth. In this case, all the descendants of the inputted node have already being processed by the RNN, and the next inputted node will necessarily be its parent or sibling.
By induction over the nodes of the tree, and using these three options we get that after applying the RNN to the root (which is the last inputted node), the only node left in the stack is the answer to the question given by the root, which is the answer to the CRQ.
We are left with showing that the stack does not overflow. We will show that the maximal number of nodes in the stack is bounded by $\textbf{mr}(T)+1$. Since the number of coordinates in the representation of each node is $d+1$, using Lemma C.2 finishes the proof. Let $v$ be some node with $\textbf{mr}(v)=k$. We will first show by induction on $k$ that at any point in time, before processing $v$, the maximal number of nodes in the stacks that are also descendants of $v$ is bounded by $k$. If $k=0$ then $v$ is a leaf a doesn’t have any children. For the induction step, assume $k>0$, and let $u_{1},u_{2}$ be the two children of $v$. If $\textbf{mr}(u_{1})=\textbf{mr}(u_{2})$, then by the definition of memory rank, they are both equal to $k-1$. Assume w.l.o.g that $u_{1}$ is processed before $u_{2}$. Note that in this case, $u_{2}$ is processed only after all the descendants of $u_{1}$ are processed. Then, by the induction step, the maximal number of nodes in the stacks that are also descendants of $u_{2}$ is bounded by $k-1$. Hence, the maximal number of nodes in the stack that are descendants of $v$ is $k$, which includes $u_{1}$. If $\textbf{mr}(u_{1})\neq\textbf{mr}(u_{2})$, assume w.l.o.g that $\textbf{mr}(u_{1})>\textbf{mr}(u_{2})$, then $u_{1}$ is processed before $u_{2}$ (resp. $u_{2}$ before $u_{1}$). In this case, by the induction case, again the number of nodes in the stack the are descendants of $v$, are either descendants of $u_{1}$, in which case no nodes that are descendants of $u_{2}$ have being processed, or descendants of $u_{2}$, in which case $u_{1}$ is in the stack, as well as at most $k-1$ descendants of $u_{2}$. This finished the induction.
Now, we will show that the number of nodes that are not descendants of $v$ and are in the stack is bounded by $1+\textbf{mr}(T)-k$. This will be by induction over $\mathcal{P}^{(i)}(v)$, namely the ancestors of $v$. We assume that $\textbf{mr}(k)=v$. If $\textbf{mr}(\mathcal{P}(v))=k$, then $v$ is processed before its sibling, which by the induction we showed before, the number of nodes in the stack that are descendant of $\mathcal{P}^{(i)}(v)$ is bounded by $k$. If $\textbf{mr}(\mathcal{P}(v))=k+1$, then the sibling of $v$ either have a memory-rank of $k$ or larger than $k$. In both cases it can be processed before $v$ (as well as all its descendants), hence the number of nodes in the stack that are descendants of $\mathcal{P}(v)$ is bounded by $k+1$. Using the same induction argument as before, applied for the sibling of $\mathcal{P}^{(i)}(v)$, we get that the number of nodes the are descendants of $v$ that are in the stack is bounded by $\textbf{mr}(\mathcal{P}^{(i+1)}(v))+1$.
Combining the two inductive arguments above, we get that when $v$ is processed into the RNN, the number of nodes in the stack is bounded by $\textbf{mr}(T)+1$. This finishes the correctness proof.
By induction on $\textbf{mr}(v)$. If $\textbf{mr}(v)=0$, then it is a leaf. By Algorithm 1, since $v$ doesn’t have children, the only node that is processed before it can be its sibling. When its sibling was processed necessarily all its children were already processed, hence there is at most $1$ node in the stack. Assume that if $\textbf{mr}(u)=k$, then there are at most $k+1$ nodes in the stack, and let $v$ with $\textbf{mr}(v)=k+1$. Denote by $u_{1}$ and $u_{2}$ the children of $v$, and assume w.l.o.g that $\textbf{mr}(u_{1})\geq\textbf{mr}(u_{2})$. If $\textbf{mr}(u_{1})>\textbf{mr}(u_{2})$, then $\textbf{mr}(v)=\textbf{mr}(u_{1})$. This means that $u_{1}$ was processed before $u_{2}$. Hence, $u_{1}$ is in the stack, and $u_{2}$ was processed right before $v$, which by induction shows that there are at most $k+1$ nodes in the stack. If $\textbf{mr}(u_{1})=\textbf{mr}(u_{2})$, the by the definition of memory-rank, it necessarily happen that they are both equal to $k$. Assume w.l.o.g that $u_{1}$ is processed before $u_{2}$ (the order of them being processed is chosen arbitrarily in this case). Hence, when processing $u_{2}$ there are at most $k$ nodes. ∎
C.3 Proof of Theorem 5.5
We begin with proving the theorem for balanced binary trees and then generalize to other trees. Let $T$ be a rooted balanced binary tree of size $n$ with a given ordering of the nodes denoted as: $v_{1},\dots,v_{n}$. Let $k:=\log_{2}(n+1)$. We define the vectors $\mathbf{t}_{0}=\begin{pmatrix}0\\ 1\end{pmatrix},~{}\mathbf{t}_{1}=\begin{pmatrix}1\\ 0\end{pmatrix},~{}\mathbf{t}_{\texttt{null}}=\begin{pmatrix}0\\ 0\end{pmatrix}$. We now define a set of CRQs given by $\mathbf{x}_{v_{i}}\in\{\mathbf{t}_{0},\mathbf{t}_{1},\mathbf{t}_{\texttt{null}}\}$ for all $i\in[n]$. We will show that given an RNN that solves all CRQs with this tree and this ordering of the nodes must have a hidden state containing $\Omega(\log(n))$ bits.
As in Lemma 4.2, an intuitive way to look at these CRQs is as Boolean formulas. For leaf nodes $\mathbf{t}_{0}$ represents 00 and $\mathbf{t}_{1}$ represents $1$. For non-leaf nodes, $\mathbf{t}_{0}$ represents $\land$ and $\mathbf{t}_{1}$ represents $\lor$. The vector $\mathbf{t}_{\texttt{null}}$ is used by a hidden node to pass values up the tree without changing them, as described below.
Let $u_{1}$ be the last leaf node in the prescribed ordering. Denote its ancestors as $u_{2},\dots,u_{k}$, where $u_{k}$ is the root. Also, for each $u_{i}$ with $i\neq 1$ denote its other child as $w_{i-1}$ (i.e. the sibling of $u_{i-1}$). Suppose that the current input to the RNN is $u_{1}$. We will show that the RNN needs at least $k$ bits in its memory at this point in the calculation to solve the CRQ.
We consider the set of CRQs defined by the following rules:
  1. 1.
    Each of the nodes $w_{1}$ and $u_{1},\dots,u_{k}$ is equal to either $\mathbf{t}_{0}$ or $\mathbf{t}_{1}$.
  2. 2.
    For each node $w_{2},\dots,w_{k-1}$, all its leaf descendants are identical to one another. Either they are all $\mathbf{t}_{0}$, or they are all $\mathbf{t}_{1}$. Its non-leaf descendants, including itself, are all equal to $\mathbf{t}_{\texttt{null}}$.
To solve the sub-question represented by $u_{2}$, the RNN must use $1$-bit of memory to remember $w_{1}$, since by assumption, $w_{1}$ precedes $u_{1}$ in the sequence. Similarly, to solve the sub-question represented by $u_{2}$, the RNN must know the answer to $w_{2}$. Because of rule 2, each subquestion in the subtree rooted at $w_{2}$ results in a tie. Because all the leaves in this subtree are identical, the answer to the subquestion rooted at $w_{2}$ equals any of its descendant leaves. These leaves all precede $u_{1}$ in the sequence, so we must use 1-bit of memory to store their value. For each node $u_{2},\dots,u_{k}$, the RNN must store at least $1$ bit of memory at the point when it processes $u_{1}$. In total, this amounts to $\log(n)-1$ bits of memory that needs to be transferred through the hidden state of the RNN.
Now, let $T$ be some rooted tree of size $n$, which is non necessarily a balanced binary tree. Denote by $\ell:=\textbf{mr}(T)$, and recall this means that the root has a memory-rank of $\ell$. We will find a subset of $T$ that forms a balanced binary tree whose depth equals $\ell$. We will construct all the other nodes to simply copy their inputs up the tree, and appeal to the case of a balanced binary tree proved above.
We first show how to find such a subset. If the root has two children with a memory rank of $\ell-1$, denote the root by $u_{1}$. Otherwise, exactly one of its children has a memory-rank of $\ell$; search recursively through its children until a node with two children of rank $\ell-1$ is found and label it $u_{1}$. We now apply the above procedure recursively on the two children of $u_{1}$. This yields two nodes, $u_{2}$ and $u_{3}$, each of which has two children with a memory rank of $\ell-2$. We apply this procedure recursively until we defined nodes $u_{1},\dots,u_{2^{\ell}}$, where each node is either a leaf or has two children with the same memory-rank as one another. By the definition of memory rank, this process will generate a subset that forms a (possibly non-consecutive) complete binary tree.
We now construct a family of CRQs on the original tree whose answer always equals that of the complete binary subset $\{u_{1},\ldots,u_{2^{\ell}}\}$. Nodes in the subset are labeled with $\mathbf{t}_{0},\mathbf{t}_{1},\mathbf{t}_{\texttt{null}}$, as above, except we append an extra 1 to each of these vectors:
$\mathbf{t}_{0}=\begin{pmatrix}0\\ 1\\ 1\end{pmatrix},\mathbf{t}_{1}=\begin{pmatrix}1\\ 0\\ 1\end{pmatrix},\mathbf{t}_{\texttt{null}}=\begin{pmatrix}0\\ 0\\ 1\end{pmatrix},$ (9)
Each node of the original tree that is not in the subset is labeled with one of the following two vectors: $\mathbf{t}_{+}=\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}$ or $\mathbf{t}_{-}=\begin{pmatrix}0\\ 0\\ -1\end{pmatrix}$. Specifically, leaf nodes that are not in the subset are labeled with $\mathbf{t}_{-}$, and non-leaf nodes are labeled with $\mathbf{t}_{+}$. To see that the answer to this CRQ equals that of the CRQ corresponding to the balanced binary subset, it suffices to show that the vectors $\mathbf{t}_{+}$ and $\mathbf{t}_{-}$ are never chosen over $\mathbf{t}_{0}$ and $\mathbf{t}_{1}$ at any subquestion. First, we show that $\mathbf{t}_{-}$ is never chosen over $\mathbf{t}_{0}$ and $\mathbf{t}_{1}$. This can be seen be enumerating all the possible values that the parent (non-leaf) node can take on ($\mathbf{t}_{0},\mathbf{t}_{1},\mathbf{t}_{\texttt{null}},\mathbf{t}_{+}$) and seeing that each prefers $\mathbf{t}_{0}$ and $\mathbf{t}_{1}$ to $\mathbf{t}_{-}$. Second, $\mathbf{t}_{+}$ is never even compared to $\mathbf{t}_{0}$ and $\mathbf{t}_{1}$, because no leaves are labeled with $\mathbf{t}_{+}$.
We now define a CRQ over $u_{1},\dots,u_{2^{\ell}}$ the same way as in the balanced binary tree case using $\mathbf{t}_{0},\mathbf{t}_{1},\mathbf{t}_{\texttt{null}}$. By construction, the depth of this synthetic tree is $\ell=\textbf{mr}(T)$, so the RNN needs memory of size $\textbf{mr}(T)$.
Finally, if an RNN solves all CRQs on trees of size $n$, it also solves them on balanced binary trees of size $n$. Thus, the number of required bits in the hidden state of the RNN is $\Omega(\log(n))$.
Appendix D Extension of the RNN solution to non-binary trees
D.1 Converting non-binary to binary trees
Suppose we have a CRQ over a tree $T$ with $n$ nodes and max degree $k$. We can transform it to an equivalent CRQ over a binary tree $T^{\prime}$ with $2n$ nodes and max degree $2$. By equivalent we mean that the answer to the two CRQs are the same.
The construction is as follows: Let $v$ be some node with degree $k$ and denote its leaves by $u_{1},\dots,u_{k}$. Suppose that $k=2^{\ell}$ for some integer $\ell$. We create a balanced binary tree with at most $2k$ nodes and depth $\ell$. The leaves will have vectors $\mathbf{x}_{u_{1}},\dots,\mathbf{x}_{u_{k}}$ and all the non-leaf nodes have a vector equal to $\mathbf{x}_{v}$. It is easy to see that the answer to the sub-question of the root of this tree is the same as the answer to the sub-question of the node $v$ from the original tree. The number of nodes is increased by a factor of at most $2$. If $k$ is not a power of $2$, we can duplicate the last leaf node until the number of leaves is a power of $2$ and use the same construction.
Doing this for every node increases the size of the tree by a factor of at most $2$, and the resulting CRQ will have the same answer as the original CRQ.
D.2 Memory-rank for non-binary trees
In this sub-section we explain how to extend some of the results from Section 5 to non-binary trees. We first define memory-rank for general trees:
Definition D.1 (Memory Rank for non-binary trees).
Let $T=(V,E)$ be a rooted tree. The memory rank of a node $v\in V$ is defined recursively as: $\textbf{mr}(v)=0$ if $v$ is a leaf. If $v$ is not a leaf, suppose it has $k$ children denoted by $u_{1},\dots,u_{k}$ where $\textbf{mr}(u_{1})\geq\dots\geq\textbf{mr}(u_{k})$, then:
$\textbf{mr}(v)$ $=\max\left(\textbf{mr}(u_{1}),\textbf{mr}(u_{2})+1,\dots,\textbf{mr}(u_{k})+k-1\right)$
It is straightforward to see that this definition generalizes Definition 5.3. Also, Algorithm 1 readily generalizes to Definition D.1, since it has no dependence on the degree of each node.
Extending Theorem 5.4 to general trees with maximal degree of $k$ can be done in a straightforward way by having an RNN with hidden dimension of $O(dk\log(n))$ and $O(k)$ layers. The idea is to add for each node an additional coordinate representing its degree. Now, in the proof of Theorem 5.4, when the last two nodes are extracted from the stack (layer $3$ in the construction), we split this into $k$ layers. Each layer checks the degree of the parent node from $2$ until $k$ and for degree $i$ extracts the last $i$ nodes from the stack. This can be implemented by a ReLU network of depth $O(k)$. Finally, the node with the highest inner product to the parent is inserted back to the stack with an updated depth (similarly to layers $4$ and $5$ from the proof of Theorem 5.4).
We believe it is also possible to extend this construction to having depth $O(1)$ and hidden dimension $O(d\log(n))$ for general trees by changing the sorting algorithm. This can be done by splitting the argmax operation over $k$ inputs into $k-1$ argmax operations, each over $2$ inputs, where these operations are nested. We leave this construction for future works.
Appendix E Proofs from Section 6
The main idea of the construction is to iteratively use the self-attention layer constructed in Theorem 4.1, while wrapping it with a construction that allows outputting CoT tokens in the right order.
By Lemma A.1 there exist vectors $\mathbf{z}_{1},\dots,\mathbf{z}_{n}\in\{\pm 1\}^{4\log(n)}$ such that $|\left\langle\mathbf{z}_{i},\mathbf{z}_{j}\right\rangle|\leq 3\log(n)$ for any $i\neq j$. We will use these vectors twice, and for two different use cases. The first one is similar to the construction of Theorem 4.1. Namely, for each node $v\in V$ that is not a leaf, we set one of the vectors $\mathbf{z}_{i}$ to be corresponded with this node and denote it as $\mathbf{z}_{v}$.
We also assume that the tree has a reverse BFS ordering. This means that we do a standard BFS ordering of the tree, and order of node $i$ in the reverse order is $n-i+1$. For a node $v$, let $I(v)$ be be its place in the reverse BFS ordering. We also correspond one of the vectors defined above to each place in this order we denote this vector as $\mathbf{w}_{I(v)}$. For the root $v_{r}$ we have $I(v_{r})=n$, and define $\mathbf{w}_{I(v_{r})+1}=\mathbf{w}_{1}$. We use a different notation for the BFS ordering to not confuse them with the other set of positional encoding, although this is the same set of vectors. It is also possible to find another set of vectors with the same property, although it won’t matter for the construction. The embedding for a leaf node $u$ and a non-leaf node $v$ are defined as:
$\mathbf{t}_{u}:=\begin{pmatrix}\mathbf{x}_{u}\\ \bm{0}_{d}\\ \mathbf{z}_{\mathcal{P}(u)}\\ \bm{0}_{4\log(n)}\\ \mathbf{w}_{I(u)}\\ \mathbf{w}_{I(u)+1}\\ 0\end{pmatrix},~{}~{}\mathbf{t}_{v}:=\begin{pmatrix}\mathbf{x}_{v}\\ \bm{0}_{d}\\ \mathbf{z}_{v}\\ \mathbf{z}_{\mathcal{P}(v)}\\ \mathbf{w}_{I(v)}\\ \mathbf{w}_{I(v)+1}\\ 1\end{pmatrix}\in\mathbb{R}^{2d+16\log(n)+1}~{}.$ (10)
As in Theorem 4.1, all the coordinates of the positional encoding are in ±$\pm\{1\}$, and the matrices that operate over these coordinates will have all their entries in $\{-1,0,1\}$. We will first give an intuitive explanation of the embedding and how the transformer will work. The embedding first contains the vector $\mathbf{x}$ corresponding to the node. It also contains its own positional encoding $\mathbf{z}_{v}$ if its a non-leaf, and the positional encoding of its parent. Next, it contains the vector representing its place in the reverse BFS ordering, as well as the subsequent vector in the order. The last coordinate is a flag for whether a node is a leaf.
The transformer will contains $2$ self-attention layers, and an MLP with $O(1)$ layers after each self-attention layer. The first self-attention layer and the MLP afterwards is similar to the construction in Theorem 4.1. Their goal is to solve the sub-questions in the CRQ. This layer will use the $\mathbf{z}$’s positional encoding vectors. The second self-attention layer will make sure that the last CoT token that was created will attend to the next token corresponding to the node after it in the reverse BFS ordering. This way, each CoT token will provide an answer to one sub-question of the CRQ, and using the reverse BFS order we make sure that any subsequent questions will have the questions of their children already solved.
We now turn to the formal construction: Let $c\geq 1$ be some constant such that $\left\lVert\mathbf{x}\right\rVert^{2}\leq c$ for every $\mathbf{x}\in\Gamma^{d}$. The matrices for the first self-attention layer are equal to:
$K=\begin{pmatrix}I_{d}&&&&\\ &\boldsymbol{0}_{d\times d}&&&\\ &&\sqrt{c}\cdot I_{4\log(n)}&&\\ &&&\boldsymbol{0}_{12\log(n)\times 12\log(n)}&\\ &&&&\sqrt{3c}\end{pmatrix}~{}$
$Q=\begin{pmatrix}I_{d}&&&&\\ &\boldsymbol{0}_{d\times d}&&&\\ &&\sqrt{c}\cdot I_{4\log(n)}&&\\ &&&\boldsymbol{0}_{12\log(n)\times 12\log(n)}&\\ &&&&-\sqrt{3c}\end{pmatrix}~{},$
$V=\begin{pmatrix}\boldsymbol{0}_{d\times d}&I_{d}&\\ I_{d}&\boldsymbol{0}_{d\times d}&\\ &&\boldsymbol{0}_{16\log(n)\times 16\log(n)+1}\end{pmatrix}~{}.$
We will first show to which each token can attend to. Let $v$ be some non-leaf node, and $u$ a child of $v$ which is a leaf, and $w$ some other node. We will show that $v$ can only attend to $u$:
$\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{u}$ $=\left\langle\mathbf{x}_{v},\mathbf{x}_{u}\right\rangle+c\left\lVert\mathbf{z}_{v}\right\rVert^{2}\geq 4c\log(n)-c$
$\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{v}$ $=\left\lVert\mathbf{x}_{v}\right\rVert^{2}+c\left\lVert\mathbf{z}_{v}\right\rVert^{2}-3c\leq 4c\log(n)-2c$
$\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{w}$ $=\left\langle\mathbf{x}_{v},\mathbf{x}_{u}\right\rangle+c\left\langle\mathbf{z}_{v},\mathbf{z}_{u}\right\rangle-3c\leq 4c\log(n)-2c$
This shows that $\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{u}\geq\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{v}$ and $\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{u}\geq\mathbf{t}_{v}^{\top}K^{\top}Q\mathbf{t}_{w}$. In particular, $v$ will attend to its child with $\max_{u\in\mathcal{C}(v)}\left\langle\mathbf{x}_{v},\mathbf{x}_{u}\right\rangle$, which is the correct answer to the sub-question represented by $v$. After applying the $V$ matrix and adding the residual connection, the to embedding of $v$ is equal to:
$\tilde{\mathbf{t}}_{v}=\begin{pmatrix}\mathbf{x}_{v}\\ \mathbf{x}_{u}\\ \mathbf{z}_{v}\\ \mathbf{z}_{\mathcal{P}(v)}\\ \mathbf{w}_{I(v)}\\ \mathbf{w}_{I(v)+1}\\ 1\end{pmatrix}~{},$
where $\mathbf{x}_{u}$ maximizes the inner product with $\mathbf{x}_{v}$ over its children. The first layer of the MLP doesn’t change the input. This can be easily done with a ReLU network since $\sigma(z)-\sigma(-z)=z$ for every $z\in\mathbb{R}$.
The second self-attention layer will have the following weights:
$K=\begin{pmatrix}\boldsymbol{0}_{2d\times 2d}&&&\\ &\boldsymbol{0}_{4\log(n)\times 4\log(n)}&I_{4\log(n)}&\\ &\boldsymbol{0}_{4\log(n)\times 4\log(n)}&\boldsymbol{0}_{4\log(n)\times 4\log(n)}&\\ &&&0\\ \end{pmatrix},$
$Q=\begin{pmatrix}\boldsymbol{0}_{2d\times 2d}&&&\\ &I_{4\log(n)}&\boldsymbol{0}_{4\log(n)\times 4\log(n)}&\\ &\boldsymbol{0}_{4\log(n)\times 4\log(n)}&\boldsymbol{0}_{4\log(n)\times 4\log(n)}&\\ &&&0\\ \end{pmatrix}~{},$
$V=\begin{pmatrix}\boldsymbol{0}_{d\times d}&I_{d}&&&\\ \boldsymbol{0}_{d\times d}&\boldsymbol{0}_{d\times d}&&&\\ &&\boldsymbol{0}_{4\log(n)\times 4\log(n)}&I_{4\log(n)}&\\ &&\boldsymbol{0}_{4\log(n)\times 4\log(n)}&\boldsymbol{0}_{4\log(n)\times 4\log(n)}&\\ &&&&I_{8d\log(n)+1}\end{pmatrix}~{}.$
The second MLP will again just copy the input, and we will not use the residual connection for the second layer  [^3] It is always possible to ignore the residual connection (if it exists) by doubling the embedding dimension. This can be done by adding zero coordinates, then use the $V$ matrix to copy the original vector intro those added coordinates, and use another linear layer from the MLP to move back the vector to the first coordinates while removing what was added from the residual connection..
We now turn to prove the correctness of the construction. The second self-attention layer forces each token $v$ to attend to the token $u$ with $I(u)=I(v)+1$, namely the consecutive token in the reverse BFS order. The $V$ matrix of the second self-attention layer transform the embedding of this token to be similar to that of an embedding of a node. Suppose we constructed $k-1$ CoT tokens already, and for a node $v$, $I(v)=k$. By the ordering, this means that all the descendants of $v$ have being processed. Also, if $v$ have descendants that are non-leaves, then the first layer correctly solved the corresponding sub-question to them, and they were outputted as CoT tokens with a similar embedding to a leaf node. Hence, in the current iteration, the token corresponding to $v$ will necessarily attend the token corresponding to its child that solves the sub-question. In addition, the last token CoT token that was created will necessarily attend to the token corresponding to $v$ by the construction of the second self-attention layer. This means that the new CoT token will be the answer to the sub-question corresponding to $v$, with and embedding similar to a leaf. This is true for any node $v$, hence the last CoT token will contain the answer to the root, which is the answer to the CRQ.
Appendix F General CRQs
In this appendix we present a generalization of CRQs by allowing different operators. For simplicity we only consider binary trees. Given a CRQ over a non-binary tree we can always define an equivalent CRQ over a binary tree, where the number of nodes increase by a factor of at most $2$, see Appendix D. We begin with a formal definition:
Definition F.1.
A General Compositional Reasoning Question (GCRQ) is a rooted binary tree $T=(V,E)$ with root $v_{r}\in V$, and a set of operators $f_{1},\dots,f_{k}:\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$. Each node $v\in V$ is labeled by a vector $\mathbf{x}_{v}\in\Gamma^{d}$, and non-leaf nodes are also labeled by a number $i_{v}\in[k]$. Here $\Gamma\subset\mathbb{R}$ is some finite vocabulary of constant size and $d\in\mathbb{N}$.
The answer to the GCRQ is defined as $\mathcal{A}(v_{r})$ where the function $\mathcal{A}$ is defined recursively: For a leaf $u\in V$ we define $\mathcal{A}(u)=\mathbf{x}_{u}$. For a non-leaf node $v\in V$ with children $u_{1},u_{2}$ we define:
$\mathcal{A}(v)=f_{i_{v}}(\mathbf{x}_{v},\mathcal{A}(u_{1}),\mathcal{A}(u_{2}))~{}.$
We refer to every node $v\in V$ which is not a leaf as a sub-question.
For a concrete example, we show how to implement arithmetic problems using GCRQs. Let $\Gamma=\{0,\pm 1,\dots,\pm A\}$ for some constant integer $A$. Also assume that $d=1$. We define the operators: $f_{+},f_{-},f_{\times},f_{\div}$ as:
$f_{+}=(\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}})=\mathbf{x}_{u_{1}}+\mathbf{x}_{u_{2}},\quad f_{-}=(\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}})=\mathbf{x}_{u_{1}}-\mathbf{x}_{u_{2}},$
$f_{\times}=(\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}})=\mathbf{x}_{u_{1}}\times\mathbf{x}_{u_{2}},\quad f_{\div}=(\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}})=\mathbf{x}_{u_{1}}\div\mathbf{x}_{u_{2}}~{}.$
Note that each such operator can be represented by a ReLU network. For $f_{+}$ and $f_{-}$ this is clear, since these are linear operators. For $f_{\times}$ we can use Lemma C.3 to approximate $(x,y)\mapsto x\cdot y$ up to an accuracy of $\frac{1}{4}$ with a $2$-layer network. Since the alphabet is finite, We can threshold over integers using another ReLU layer to get an exact representation of multiplication over a finite set of integers. A similar construction can be used for $f_{\div}$.
We now explain how to change our constructions to work with GCRQs. Note that the only place in the proofs of Theorem 4.1 and Theorem 6.1 we use the fact that the operator is argmax, is where we use a single self-attention layer to calculate inner products. Instead, we can use such a layer so that the parent node will attend to both its children with the output being a concatenation of their vectors. In more details, assume that each two siblings $u_{1}$ and $u_{2}$ in the tree are labeled by vectors $\begin{pmatrix}\mathbf{x}_{u_{1}}\\ \boldsymbol{0}_{d}\end{pmatrix}$ and $\begin{pmatrix}\boldsymbol{0}_{d}\\ \mathbf{x}_{u_{2}}\end{pmatrix}$. Then, their parent can attend to both of them with the same weight using the positional embeddings, in the same way as done in the proofs of the mentioned theorems. We also construct the MLP so that its input will include the reference to the operator $i_{v}$. The MLP will simulate an ”if“ statement over the $k$ operators and apply the relevant one over the vectors $\mathbf{x}_{v},\mathbf{x}_{u_{1}},\mathbf{x}_{u_{2}}$. This ”if“ statement can be simulated by a $k$-layer ReLU network (we assume the $k$ is independent of $n$).
A similar construction will work for the proof of Theorem 5.4, where here the MLP will output the last two inputs from the stack and apply the operator $i_{v}$. Again, it will increase the depth of the MLP by $k$. We leave exact constructions for future work.